Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2007-11-23
ISS On-Orbit Status 11/23/07

FE-1 Yuri Malenchenko performed Part 2 of the ground-controlled test of the Russian data telemetry system's MKO multiplex exchange channel, using BSR-TM payload data telemetry (TM) and the 4PrNP-6 data gathering application of the BITS2-12 Onboard Telemetry Measurement System.

(The test, controlled from the RSS2 laptop, began 11/21 with Yuri switching from the regular 128-byte TM frame to a 206-byte format, for TsUP to run tests from RGS (Russian Ground Sites). Today, in Part 2 as per plan the FE-1 returned the BSR-TM to the nominal 128-byte format.)

Malenchenko also transferred measurements & imagery from the ESA/RSC-Energia experiment ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS) to OCA for subsequent downlink to the ground, after yesterday's first repositioning of the spectrometer. (ALTCRISS uses the AST spectrometer to monitor space radiation in the Russian segment (RS).)

FE-2 Dan Tani started his workday with another 60-min of personal time for general orientation (station familiarization & acclimatization), as is standard daily rule for the first two weeks after starting station residency.

The crewmembers performed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the TVIS treadmill (CDR, FE-1, FE-2), RED resistive exerciser (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

Later, the CDR transferred the crew's exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

After the physical workout, CDR Whitson took air samples for the periodic (weekly) atmospheric status check for ppO2 (Partial Pressure Oxygen) and ppCO2 (pp Carbon Dioxide), using the hand-held CSA-O2 (Compound Specific Analyzer-Oxygen sensor) and CDMK (CO2 Monitoring Kit). Batteries were to be replaced if necessary. (Purpose of the 10-min. activity is to trend with MCA (Major Constituents Analyzer), i.e., to correlate the hand-held readings with MCA measurements.)

Dan Tani filled out the regular FFQ (Food Frequency Questionnaire), his third, on the MEC (Medical Equipment Computer). (By means of these FFQs, U.S. astronauts keep a personalized log of their nutritional intake over time on special MEC software. Recorded are the amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins. At TsUP/Moscow, food specialists are currently preparing the Russian food 'menu' for delivery by Progress M-63/28P next February. 28P will carry 'bonus food' for Peggy and Yuri, plus about 15 kg of fresh food items (apples, grapefruit, oranges, lemons, garlic) in two containers.)

The FE-1 collected the periodic water samples downstream of the BKO multifiltration/purification column unit, to be returned to the ground on STS-122/1E for monitoring the quality of the water being fed to the Elektron O2 generator by the BKO. (The SRV-K2M Condensate Water Processor's BKO column removes dissolved mineral and organic impurities from the condensate.)

Malenchenko also completed the daily routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM (Service Module), including ASU toilet facilities systems/replaceables.

Later, Yuri conducted the daily IMS (Inventory Management System) maintenance, updating/editing its standard 'delta file' including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur),- again a voluntary task from his 'time permitting' discretionary task list.

Working off the voluntary U.S. 'job jar' task list, the CDR performed computer maintenance, doing the regular bi-weekly reboots of the SSC (Station Support Computer) File Server and OCA Comm Router laptops.

Spending time in the Airlock (A/L) for final preparations for tomorrow's EVA-12 'Charlie', Peggy Whitson and Dan Tani -

Configured the DCS-760 digital camera to be taken outside,
Recharged one DCS battery for the EMU Prebreathe activity in the A/L (when off station power),
Prepared the A/L EL (Airlock Equipment Lock) for the Campout & spacewalk,
Tagged up with ground specialists at ~10:40am EST for reviewing EVA particulars, and
Underwent the standard pre-EVA PMC (Private Medical Conference) via S- & Ku-band audio/video.
Also in the A/L, Peggy terminated the regeneration of METOX (Metal Oxide) canisters #0020 & #0021 in the bakeout oven. (The reusable METOX CO2 absorbers will not be used in the EMUs for EVA-12, but in the A/L during tonight's 'Campout' lockout. During the spacewalk, CO2 in the EMUs will be scrubbed with non-reusable LiOH (Lithium Hydroxide) canisters, with more 'lifetime' than METOX cans, to protect against a possible timeline extension of EVA 'Charlie' that - like EVA 'Bravo' - is considered a "full" EVA. This means that all timeline content needs to be done in one EVA due to Node-2 Thermal Loop System shutdown timing. Also, working with ammonia (NH3) fluid couplings could require additional time for decontamination procedures.)

Before the Campout, the CDR powered down onboard amateur (ham) radio equipment (Kenwood in SM, Ericsson in FGB) to prevent RF (radio frequency) interference with the EMUs.

Yuri again prepared three Russian 'Pille-MKS' radiation dosimeters, recorded their dosages and equipped each of the two EMUs (Extravehicular Mobility Unit) with a radiation sensor (A0309/CDR & A0310/FE-2). (A third sensor, A0308, was placed in the SM on the PULT reader for background readings.)

At ~3:55am EST, Yuri linked up with TsUP stowage specialists via S-band to conduct the weekly IMS tagup, discussing stowage issues and equipment locations. (Issues of discussion today included the search for a 'lost' cable of the Russian Agat-2 Audio & Video Complex, needed for the future.)

At ~1:45pm, the crew is scheduled for their third weekly tagup with the Lead Flight Director at JSC/MCC-H via S-band/audio (S/G 2 (Space-to-Ground 2) phone patch via SSC-10 (Station Support Computer 10)).

At ~2:50pm EST, Whitson & Tani will begin their Campout mask prebreathe, lasting until ~4:00pm and followed by hatch closure by Yuri and overnight lockout in the A/L at 10.2 psi for denitrogenation.

Before his sleep time, Malenchenko will again isolate the Node-2 for the EVA (after its ground-commanded deactivation at ~12:00pm), by uncapping Harmony's Stbd & Port PPRVs (Positive Pressure Relief Valves), temporarily stowing the caps, closing the Node-2 Port IMV (Intermodular Ventilation) valve manually and closing Node-2 aft hatch.

EVA-12 'Charlie' Preview: After an 8.5-hr Campout sleep period on 10.2 psi, the A/L CL (Crewlock) hatch will be cracked for the usual hygiene break/with mask prebreathe. Afterwards, the hatch will be closed again by Yuri for EVA preparations at 10.2 psi, followed by EMU purge & prebreathe, then CL depressurization and EV1/EV2 egress. Tomorrow morning, EVA-12 begins nominally at ~5:55am EST, probably much earlier in real-time (knowing Peggy), with a 6h 30m timeline duration.

Dedicated to Part 2 of Node-2 outfitting (i.e., connecting & configuring the second half of Harmony's fluid, power and cooling jumpers), its objectives are to:

Remove the S0 truss NH3 (ammonia) shunt jumper on Port;
Configure vent tools;
Remove Node-2 fluid QD (quick disconnect) caps;
Vent & stow the Port NH3 shunt jumper;
Relocate an APFR (Articulating Portable Foot Restraint) from Lab WIF-11 (Worksite Interface Fixture 11) to Lab WIF-12;
Relocate Node-2 Loop B fluid tray to the Lab & attach it;
Deploy Node-2 Loop B fluid tray hinged section;
Mate S0 fluid QDs, then open S0 valves and 2 fluid QDs;
Connect two Node-2 fluid line heater cables;
Connect five Node-2 Stbd avionics cables to Node-2;
Release Node-2 Stbd CBM (Common Berthing Mechanism) petal launch locks (8 latches);
Mate remaining SSPTS (Station-to-Shuttle Power Transfer System) umbilical (connector J16A) to PMA-2 (ctr P16);
Install Lab/Node-2 gap spanner on two handrails (Lab & Node 2); and
Remove Stbd SARJ (Solar Alpha Rotary Joint) Cover 7, using PGT (Pistol Grip Tool) on its 6 bolds.)
Get-ahead tasks are:

Release Node-2 Port CBM petal launch locks (8 latches). (Nadir CBM launch locks remain closed at this time to prevent the unlatched petals from opening (due to lost pin);
Re-installation of CETA (Crew & Equipment Translation Aid) Light Fixture;
Clean-up of Lab MMOD (micrometeoroid/orbital debris) shield.
No CEO (Crew Earth Observation) photo targets uplinked for today.

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use